i dont understand how to write an equation for the line that is parallel to y=(4/3)x-7 ; and contains given point, (7,2).
Found 3 solutions by rfer, mananth, stanbon:
Answer by rfer(16322)

solve for b
y=mx+b
2=4/3*7+b
2=9.3+b
-7.3=b
------------
y=4/3x-7.3

Answer by mananth(16949)

y = 4/3 x -7
Compare this equation with y=mx+b
slope m = 4/3
The slope of a line parallel to the above line will be the same m= 1.33 ,point ( 7 , 2 )
Find b by plugging the values of m & the point in
y=mx+b
2 = 9.33 + b
b= -7.33
m= 4/3
Plug value of the slope and b in y = mx +b
The required equation is y = 4/3x -22/3
y = 4/ 3 x-22/ 3
m.ananth@hotmail.ca

Answer by stanbon(75887)

i dont understand how to write an equation for the line that is parallel to
y=(4/3)x-7 ; and contains given point, (7,2).
-----
The given line has a slope of m = 4/3
----
Any line parallel to it must have slope = 4/3.
--------
Form of the Equation:
y = mx + b
----
You know y = 2 when x = 7 and the slope is 4/3.
Solve for "b":
2 = (4/3)*7 + b
2 = 38/3 + b
b = 6/3-38/3
---
b = -32/3
----
Equation:
y = (4/3)x -(32/3)
========================
Cheers,
Stan H.
=================